OpenStudy (legomyego180):

When Integrating by Parts multiple times, do you carry the negative in front of the integral in the forumla uv-(integral)vdu?

OpenStudy (legomyego180):

I was asked to integrate: \[\int\limits_{}^{}x^2\cos(x)\] for context

OpenStudy (faiqraees):

The negative sign has to be multiplied with the resultant integral \[\large\rm ∫ab'=ab-(∫a'b)\ \]

OpenStudy (faiqraees):

Consult Example 9 on this page http://tutorial.math.lamar.edu/Classes/CalcII/IntegrationByParts.aspx They have a very beautiful trick for long integration by part questions

OpenStudy (legomyego180):

Thats what I thought too

OpenStudy (legomyego180):

Ah that is a nice trick. Cool the things you can do if you can recognize a pattern in math

OpenStudy (faiqraees):

Math is indeed beautiful.